OpenStudy (anonymous):

solve the equation 3r^2+2r-8=0

OpenStudy (anonymous):

\[3(r^2+(2/3)r)-8=0\]

OpenStudy (anonymous):

\[3(r^2+(3/2)r/2)-8=0\]

OpenStudy (anonymous):

\[3(r^2+3/2r+(3/4)^2)-8=0\]

OpenStudy (anonymous):

\[3(r^2+3/2r+9/16)-27/16-8=0\]

OpenStudy (anonymous):

\[3(r+3/4)^2-155/16=0\]

OpenStudy (anonymous):

\[3(r+3/4)^2=155/16\]

OpenStudy (anonymous):

\[(r+3/4)^2=\frac{ 155 }{ 16 }/3\]

OpenStudy (anonymous):

\[(r+3/4)^2=155/48\]

OpenStudy (anonymous):

\[r+3/4=\sqrt{(\frac{ 155 }{ 48 })}\]

OpenStudy (anonymous):

\[r=\pm \sqrt{465}/12-3/4\]